ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 88, 1994

Abhyankar, S. S. Nice equations for nice groups  1

Adams, S. Representation varieties of arithmetic groups and polynomial periodicity of Betti numbers

73

Adams, S. See P. Sarnak

Anderson, G. D., M. K. Vamanamurthy and M. Vuorinen Functional inequalities for the Teichmüller ring capacity

367

Binkowska, M. and B. Kamiński Entropy increase for ℤ^{d}-actions on a Lebesgue space

307

Birman, J. S. See B. Wajnryb

Blasius, D. On multiplicities for SL(n)  237

Cohen, M., S. Raianu and S. Westreich Semiinvariants for Hopf algebra actions

279

Eklof, P. C., A. H. Mekler and S. Shelah Hereditarily separable groups and monochromatic uniformization

213

Gil-Medrano, O. and P. W. Michor Geodesics on spaces of almost hermitian structures

319

Kaliman, S. Exotic analytic structures and Eisenman intrinsic measures

411

Kamiński, B. See M. Binkowska

Klein, A. A. The sum of nil one-sided ideals of bounded index of a ring

25

Larsen, M. On the conjugacy of element-conjugate homomorphisms 253
Le Merdy, C. Analytic factorizations and completely bounded maps 381
Levin, G. M. On the complement of the Mandelbrot set  189

Lin, M. and R. Wittmann Convergence of representation averages and of convolution powers

125

Madras, N. and D. Tanny Oscillating random walk with a moving boundary

333

Mekler, A. H. See P. C. Eklof

Michor, P. W. See O. Gil-Medrano

Raianu, S. See M. Cohen

Rudnick, Z. See P. Sarnak Sarnak, P., S. Adams and Z. Rudnick Betti numbers of congruence

groups

31

Shelah, S. How special are Cohen and random forcings, i.e. Boolean algebras of the family of subsets of reals modulo meagre or null

159

Shelah, S. See P. C. Eklof Tanny, D. See N. Madras

Troyanski, S. L. On some generalizations of denting points  175

Vamanamurthy, M. K. See G. D. Anderson Vuorinen, M. See G. D. Anderson Wajnryb, B. and J. S. Birman Errata: Presentations of the mapping

class group

425

Westreich, S. See M. Cohen Wittmann, R. See M. Lin